OpenStudy (anonymous):

Help for MEDAL : question about sequences n 1 2 3 4 5 6 f(n) 1 2 3 5 8 13

OpenStudy (anonymous):

Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2